# A Fourth-Order Cut-cell Multigrid Method

### v1

@@ 1 Introduction
    To represent an irregular domain $\Omega$ with arbitrary geometry and topology, can we have a simple and efficient scheme that is always fourth-order accurate? Furthermore, for a given threshold $\epsilon\in(0,1)$ and a Cartesian grid of size $h$, can we partition $\Omega$ into a set of cut cells whose volumes are no less than $\epsilon h^{2}$?
@@ 2 Modeling continua with Yin sets
(3) $\mathcal{Y}=\cup^{\perp\perp}_{j}{\cal Y}_{i}=\cup^{\perp\perp}_{j}\cap_{i}\operatorname{int}(\gamma_{j,i}),$
where ${\cal Y}_{j}$ is the $j$th connected component of $\mathcal{Y}$, the binary operation $\cup^{\perp\perp}$ the regularized union defined as $\mathcal{S}\cup^{\perp\perp}\mathcal{T}:=\left(\mathcal{S}\cup\mathcal{T}\right)^{\perp\perp}$, $\{\gamma_{j,i}\}$ the set of pairwise almost disjoint Jordan curves satisfying $\partial{\cal Y}_{i}=\cup_{j}\gamma_{j,i}$, and $\operatorname{int}(\gamma_{j,i})$ the complement of $\gamma_{j,i}$ that always lies at the left of an observer who traverses $\gamma_{j,i}$ according to its orientation.
@@ 3.2 Resolving the small-volume problem by cell merging
Figure 1: An illustration of Algorithm 1 in generating $\mathsf{C}_{\epsilon}^{h}(\Omega)$ in (9) with $\epsilon=0.2$ by cutting and merging cells for the domain $\Omega$. The cut cells ${\cal C}_{\mathbf{l}}$, ${\cal C}_{\mathbf{p}}$, and ${\cal C}_{\mathbf{q}}$ are regular, empty, and irregular, respectively. The symbol $\leftrightarrow$” indicates cell merging. Originally, ${\cal C}_{\mathbf{k}}\in\mathsf{C}_{\Omega}$ is an irregular cell with two small connected components, which are merged to $\mathcal{C}_{\mathbf{j}}$ at line 5 and $\mathcal{C}_{\mathbf{i}}$ at line 9, respectively. Then ${\cal C}_{\mathbf{k}}$ is removed from $\mathsf{C}_{\epsilon}^{h}(\Omega)$ and its type changed from “irregular” to “empty.” The type of ${\cal C}_{\mathbf{i}}$ is changed from “regular” to “irregular.”
@@ 4.1 Discretizing (1) on SFV cells
The case $b=0$ in (15) follows directly from (12) and the following symmetric finite-volume discretization of the gradient operator and the Laplacian:
@@
(16b) $\left\langle\Delta\phi\right\rangle_{\mathbf{i}}=\frac{1}{12h^{2}}\sum_{d}\Big(-\left\langle\phi\right\rangle_{\mathbf{i}-2\bm{\mathrm{e}}^{d}}+16\left\langle\phi\right\rangle_{\mathbf{i}-\bm{\mathrm{e}}^{d}}-30\left\langle\phi\right\rangle_{\mathbf{i}}+16\left\langle\phi\right\rangle_{\mathbf{i}+\bm{\mathrm{e}}^{d}}-\left\langle\phi\right\rangle_{\mathbf{i}+2\bm{\mathrm{e}}^{d}}\Big)+\operatorname{O}(h^{4});$
@@ 4.2.1 Poised lattice generation (PLG)
In Definition 4.3, $\mathbb{Z}^{{\scriptsize\textsf{D}}}$ captures the rectangular structure of FD grids while $K$ reflects the physics of the spatial operator being discretized. For example, to discretize the Laplacian operator, the isotropy of diffusion naturally yields the centering of $K$ at $\mathbf{q}$, whereas, for an advection operator, we set $K$ to be a lopsided box with respect to $\mathbf{q}$ because most information comes from the upwind direction; see [41, Fig. 5].
@@
[width=0.4]tikz/Stencil
Figure 3: An example of the stencil for multivariate polynomial fitting in the FV formulation for ${\scriptsize\textsf{D}}=2$ and $n=4$. The starting point $\mathbf{q}=\mathbf{i}$ is marked by $\bullet$,” the multiindices of shaded cells constitute $\mathcal{S}_{\mathrm{PLG}}$ in (19), and the thick curve segment represents the cut boundary $\mathcal{B}_{\mathbf{i}}$.
@@
the goal is to determine a vector of coefficients $\bm{\beta}=\left[\beta_{1},\cdots,\beta_{N},\beta_{\mathrm{b}}\right]^{\top}$ such that the linear combination $\bm{\beta}^{\top}\overline{\mathbf{u}}$ is an (n1)(n-1)th-order approximation of $\left\langle{\cal L}u\right\rangle$,
(21) $\forall u\in{\cal C}^{n+1}(\mathbb{R}^{\scriptsize\textsf{D}},\mathbb{R}),\quad\bm{\beta}^{\top}\overline{\mathbf{u}}=\left\langle\mathcal{L}u\right\rangle_{\mathbf{i}}+\operatorname{O}(h^{n-1}).$
@@
for a regular PLG cell ${\cal C}_{\mathbf{i}}$, the last column of $M$ in (24) is dropped.
We calculate the integrals on regular cells by six-order recursive Gauss formulas. In contrast, the integral of a scalar function over an irregular cut cell is first converted by Green’s theorem to another integral on the boundary Jordan curve and then approximated by sixth-order Gauss formulas; see [36] for more details. Together with the explicit approximation of the boundary Jodan curve with cubic splines, this integral formulation makes our method robust in that its fourth accuracy holds even at the presence of kinks on the domain boundary.
@@ 5.1 An optimal LU factorization of $A_{22}$ in (30)
where $\mathbf{0}$ represents a block matrix whose elements are all zero; the bandwidth of $A_{22,c}$ is at most 17, so is the dimension of the square blocks $A_{22,u}$ and $A_{22,l}$.
@@
By $n=4$ and Definition 4.1, an SFV cell is at the center of a 5-by-5 box of regular cells. The distance from the center of a PLG cell to $\gamma$ is at most $\frac{5\sqrt{2}}{2}h$. In the worst case, the set $K$ of feasible nodes in Definition 4.2 is a 5-by-5 box, of which the starting point $\mathbf{q}$ is a corner. Then by the total ordering in Definition 5.1, the maximum difference between the numbering of $\mathbf{q}$ and that of any multi-index in the box is $5\times\frac{5\sqrt{2}}{2}\approx 17.7$.
The above argument does not hold for a PLG cell close to the point $\gamma(0)$. However, the effects of such a PLG cell can be assimilated either into $A_{22,l}$ or into $A_{22,u}$, whose dimensions, by similar arguments as above, are at most 17.
@@ 5.2 A block smoother
The block smoother will also be applied in Algorithm 2 to the residual equation $A\mathbf{e}=\mathbf{r}$. Then the RHS residual vector $\mathbf{r}$ must be updated after each iteration.
@@ 5.3 A cut-cell V-cycle
Side-effect: $\mathbf{u}^{h}$ is updated to $(A^{h})^{-1}\mathbf{b}^{h}$.
@@
Table 1: Values of $\rho(TG)$, the spectral radius of $TG$ in (41) with $\omega=\frac{1}{2}$, for elliptic problems on various domains as specified in Section 6. In particular, the elliptic equation solved in Figure 6(b) has a cross-derivative term and the irregular boundary in Figure 8(a) is equipped with a Neumann boundary condition. For each case, we select three successively refined grids so that the most significant digits of the calculated spectral radii are the same on the two finest grids. The pairs of integers in the first row are values of $(\nu_{1},\nu_{2})$.
@@ 5.4 A cut-cell FMG cycle
For both the cut-cell V-cycle and a classical multigrid V-cycle, the convergence factor $\rho$ is a constant independent of the grid size $h$. While $\rho<1$ guarantees convergence, the solution of $A^{h}\mathbf{u}^{h}=\mathbf{b}^{h}$ entails $\operatorname{O}(\log(h^{-1}))$ V-cycles. It is well known in multigrid literature [4, p. 77–78] that this suboptimal complexity of $\operatorname{O}(\log(h^{-1}))$ V-cycles can be improved to the optimal complexity of $\operatorname{O}(1)$ FMG cycles.
@@
Output: The solution to the algebraic equation $A^{h}\mathbf{u}^{h}=\mathbf{b}^{h}$.
@@
2:  return $\text{{BottomSolver}}(A^{h},\mathbf{b}^{h})$
@@
8:return $\mathbf{u}^{h}$
Our cut-cell FMG cycle is formalized in Algorithm 3 and illustrated in Figure 5. To solve the linear system $A^{h_{f}}\mathbf{u}^{h_{f}}=\mathbf{b}^{h_{f}}$ on the hierarchy in (38), we simply invoke FMG($A^{h_{f}},\mathbf{b}^{h_{f}},\nu_{1},\nu_{2}$). Numerical tests in Section 6 show that, for $(\nu_{1},\nu_{2})=(3,3)$, each iteration of this cut-cell FMG cycle reduces the residual by a factor between 7.5 and 10.
@@ 6.3 A square minus four disks
$L^{1}$ 1.42e-03 -0.58 2.13e-03 0.62 1.38e-03 1.03 6.75e-04
$L^{2}$ 2.72e-03 -0.24 3.20e-03 0.67 2.02e-03 0.96 1.04e-03
@@ 6.4 A panda
To showcase the capability of the proposed cut-cell method in handling complex topology and geometry, we numerically solve (1) on the domain of a panda shown in Figure 9(a), which is adapted from that in [40, Figure 10] with more breakpoints so that the representation error of the complex geometry is dominated by the solution error of our cut-cell method. The complex topology and geometry of the panda pose significant challenges to a numerical solver.
@@
Despite of its $O(h^{-2})$ complexity, the generation of cut cells is very much dominated by the determination of the linear system (30), which holds even on the finest grid. Consequently, the cost of the entire initial setup displays a roughly linear growth as the grid size $h$ is reduced. Being the most expensive component of V-cycles and FMG cycles, block smoothing consumes more CPU time than the initial setup on the finest grid, since the $O(h^{-2})$ growth of its cost is higher than the linear growth.
@@ 6.5 FMG efficiency
The panda is a good representative of complex domains while the rotated square in Figure 6(b) that of the other extreme of irregular but simple domains. For the rotated square in Figure 6(b) with $h=\frac{1}{256},\frac{1}{512},\frac{1}{1024}$, we record computational costs (not shown) of the main components. The cost of generating cut cells grows quadratically with respect to the reduction of $h$ while those of all other steps in the initial setup grows linearly; the $O(h^{-2})$ complexity of FMG cycles are also confirmed. In addition, the consumed CPU time in seconds for $h=\frac{1}{1024}$ is 1.89, 2.61, 3.57, and 5.31 for the determination of (30), the entire initial setup, the block smoothing, and all FMG cycles, respectively. The cost ratio of block smoothing over all FMG cycles is $\frac{3.57}{5.31}\approx 0.67$ for the rotated square, which is very close to that ($\frac{2.38}{3.63}\approx 0.66$) for the panda, cf. Table 6. Due to the simple geometry of the rotated square, the cost ratio of the entire setup over all FMG cycles, $\frac{2.62}{5.31}\approx 0.49$, is much smaller than that ($\frac{4.55}{3.63}\approx 1.25$) for the panda test. We sum up main conclusions of the above discussions as follows.
@@
  • Block smoothing consumes two thirds of the entire CPU time of FMG cycles and is thus the most expensive component of the proposed cut-cell method.
(a) results for Section 6.1 with $h=\frac{1}{512}$
(b) results for Section 6.4 with $h=\frac{1}{2048}$
Figure 10: Performance of FMG cycles of our method with $\nu_{1}=\nu_{2}=3$ for tests on the unit square $(0,1)^{2}$, the rotated square $\Omega_{r}$, and the panda in Sections 6.1 and 6.4. The ordinates are the relative residuals/errors in $L^{\infty}$ norm, the abscissa is the iteration number of FMG cycles. The condition number of an FMG cycle is indicated by the smallest solution error that remains constant in spite of more multigrid iterations.

### v2

@@ 1 Introduction
    To represent an irregular domain $\Omega$ with arbitrary geometry and topology, can we have a simple and efficient scheme that is always fourth-order accurate? Furthermore, for a given threshold $\epsilon\in(0,1)$ and a Cartesian grid of size $h$, can we partition $\Omega$ into a set of cut cells whose volumes are between $\epsilon h^{2}$ and $2h^{2}$?
@@ 2 Modeling continua with Yin sets
(3) $\mathcal{Y}=\cup^{\perp\perp}_{j}{\cal Y}_{j}=\cup^{\perp\perp}_{j}\cap_{i}\operatorname{int}(\gamma_{j,i}),$
where ${\cal Y}_{j}$ is the $j$th connected component of $\mathcal{Y}$, the binary operation $\cup^{\perp\perp}$ the regularized union defined as $\mathcal{S}\cup^{\perp\perp}\mathcal{T}:=\left(\mathcal{S}\cup\mathcal{T}\right)^{\perp\perp}$, $\{\gamma_{j,i}\}$ the set of pairwise almost disjoint Jordan curves satisfying $\partial{\cal Y}_{j}=\cup_{i}\gamma_{j,i}$, and $\operatorname{int}(\gamma_{j,i})$ the complement of $\gamma_{j,i}$ that always lies at the left of an observer who traverses $\gamma_{j,i}$ according to its orientation.
@@ 3.2 Resolving the small-volume problem by cell merging
Figure 1: An illustration of Algorithm 1 in generating $\mathsf{C}_{\epsilon}^{h}(\Omega)$ in (9) with $\epsilon=0.2$ by cutting and merging cells for the domain $\Omega$. The cut cells ${\cal C}_{\mathbf{l}}$, ${\cal C}_{\mathbf{p}}$, and ${\cal C}_{\mathbf{q}}$ are regular, empty, and irregular, respectively. The symbol $\leftrightarrow$” indicates cell merging. Originally, ${\cal C}_{\mathbf{k}}\in\mathsf{C}_{\Omega}$ is an irregular cell with two small connected components, which are merged to $\mathcal{C}_{\mathbf{j}}$ at line 5 and $\mathcal{C}_{\mathbf{i}}$ at line 9, respectively. Then ${\cal C}_{\mathbf{k}}$ is removed from $\mathsf{C}_{\epsilon}^{h}(\Omega)$ and its type changed from “irregular” to “empty.” The type of ${\cal C}_{\mathbf{i}}$ is changed from “regular” to “irregular” whereas the type of ${\cal C}_{\mathbf{j}}$ remains “irregular.”
@@ 4.1 Discretizing (1) on SFV cells
The case $b=0$ in (15) follows directly from (12) and the following symmetric finite-volume discretization of the first and second derivatives:
@@
(16b) $\left\langle\frac{\partial^{2}\phi}{\partial x_{d}^{2}}\right\rangle_{\mathbf{i}}=\frac{1}{12h^{2}}\Big(-\left\langle\phi\right\rangle_{\mathbf{i}-2\bm{\mathrm{e}}^{d}}+16\left\langle\phi\right\rangle_{\mathbf{i}-\bm{\mathrm{e}}^{d}}-30\left\langle\phi\right\rangle_{\mathbf{i}}+16\left\langle\phi\right\rangle_{\mathbf{i}+\bm{\mathrm{e}}^{d}}-\left\langle\phi\right\rangle_{\mathbf{i}+2\bm{\mathrm{e}}^{d}}\Big)+\operatorname{O}(h^{4});$
@@ 4.2.1 Poised lattice generation (PLG)
[…]
@@
[width=0.36]tikz/Stencil
Figure 3: An example of the stencil for multivariate polynomial fitting in the FV formulation for ${\scriptsize\textsf{D}}=2$ and $n=4$. $\bullet$” marks the starting point $\mathbf{q}=\mathbf{i}$, the cells with dark shades constitute $\mathcal{S}_{\mathrm{PLG}}$ in (19), and the thick curve segment represents the cut boundary $\mathcal{B}_{\mathbf{i}}$.
@@
the goal is to determine a vector of coefficients $\bm{\beta}=\left[\beta_{1},\cdots,\beta_{N},\beta_{\mathrm{b}}\right]^{\top}$ such that the linear combination $\bm{\beta}^{\top}\overline{\mathbf{u}}$ is an (n1)(n-1)th-order approximation of $\left\langle{\cal L}u\right\rangle_{\mathbf{i}}$,
(21) $\forall u\in{\cal C}^{n+1}(\mathbb{R}^{\scriptsize\textsf{D}},\mathbb{R}),\quad\bm{\beta}^{\top}\overline{\mathbf{u}}=\left\langle\mathcal{L}u\right\rangle_{\mathbf{i}}+\operatorname{O}(h^{n-1}),$
[…]
@@
For a regular PLG cell ${\cal C}_{\mathbf{i}}$, the last column of $M$ in (24) is dropped, so are the last elements of $\bm{\beta}$ and $\overline{\bm{\phi}}$.
We calculate the integrals on regular cells by six-order recursive Gauss formulas. In contrast, the integral of a scalar function over an irregular cut cell is first converted by Green’s theorem to another integral on the boundary Jordan curve and then approximated by sixth-order Gauss formulas; see [36] for more details. Together with the explicit approximation of the boundary Jordan curve with cubic splines, this integral formulation makes our method robust in that its fourth accuracy holds even at the presence of kinks on the domain boundary.
@@ 5.1 An optimal LU factorization of $A_{22}$ in (30)
where $\mathbf{0}$ represents a block matrix whose elements are all zero; the bandwidth of $A_{22,c}$ is at most 17, so are the dimensions of the square blocks $A_{22,u}$ and $A_{22,l}$.
@@
By $n=4$ and Definition 4.1, an SFV cell is at the center of a 5-by-5 box of regular cells. A nonempty cut-cell is either an SFV cell or a PLG cell and each row of $A_{22}$ corresponds to a PLG cell. Therefore, the distance from the center of a PLG cell to $\gamma$ is at most $\frac{5\sqrt{2}}{2}h$. In the worst case, the set $K$ of feasible nodes in Definition 4.2 is a 5-by-5 box, of which the starting point $\mathbf{q}$ is at the box corner. By the total ordering in Definition 5.1, the difference between the numbering of $\mathbf{q}$ and that of any multi-index in the box is bounded by $5\times\frac{5\sqrt{2}}{2}\approx 17.7$.
The above argument does not hold in a local neighborhood of $\gamma(0)$, where the difference of the numbering of two such PLG cells might be close to the total number of PLG cells. However, the number of these pairs of PLG cells is $O(1)$ and these large differences in PLG cell numbering can be assimilated either into $A_{22,l}$ or into $A_{22,u}$, whose dimensions, by similar arguments as above, are at most 17.
@@ 5.2 A block smoother
The block smoother will also be applied in Algorithms 2 and 3 to the residual equation $A\mathbf{e}=\mathbf{r}$. Then the residual vector $\mathbf{r}$ must be updated after each iteration.
@@ 5.3 A cut-cell V-cycle
Side-effect: $\mathbf{u}^{h}$ is updated as a better approximation to $(A^{h})^{-1}\mathbf{b}^{h}$.
@@
Table 1: Values of $\rho(TG)$, the spectral radius of $TG$ in (41) with $\omega=\frac{1}{2}$, for elliptic problems on various domains as specified in Section 6. In particular, the elliptic equation solved on the rotated square in Figure 6(b) has a cross-derivative term and the irregular boundary in Figure 8(a) is equipped with a Neumann boundary condition. For each case, we select three successively refined grids so that the most significant digits of the calculated spectral radii are the same on the two finest grids. The pairs of integers in the first row are values of $(\nu_{1},\nu_{2})$.
@@ 5.4 A cut-cell FMG cycle
The convergence factor $\rho$ of a V-cycle is usually independent of the grid size $h$ and is less than 1, and thus it takes $\operatorname{O}(\log(h^{-1}))$ V-cycles to solve the linear system $A^{h}\mathbf{u}^{h}=\mathbf{b}^{h}$. It is well known from the multigrid literature [4, p. 77–78] that this suboptimal complexity of $\operatorname{O}(\log(h^{-1}))$ V-cycles can be improved to the optimal complexity of $\operatorname{O}(1)$ FMG cycles.
@@
Input: $(A^{h},\mathbf{r}^{h})$: a residual equation corresponding to the linear system (29);
@@
Output: An approximation to $(A^{h})^{-1}\mathbf{r}^{h}$.
@@
2:  return $\text{{BottomSolver}}(A^{h},\mathbf{r}^{h})$
@@
8:return $\mathbf{e}^{h}$
[…]
@@ 6.3 A square minus four disks
4th-order EB method with kink mollification; a Dirichlet condition on $\partial\Omega_{d}$
@@
$L^{1}$ 1.42e-03 0.58-0.58 2.13e-03 0.62 1.38e-03 1.03 6.75e-04
$L^{2}$ 2.72e-03 0.24-0.24 3.20e-03 0.67 2.02e-03 0.96 1.04e-03
4th-order EB method with kink mollification; a Neumann condition on $\partial\Omega_{d}$
@@ 6.4 A panda
To showcase the capability of the proposed cut-cell method in handling complex topology and geometry, we numerically solve (1) on the domain of a panda shown in Figure 9(a), which is adapted from that in [40, Figure 10] with a sufficent number of breakpoints. The same spline representation of the panda boundary is used for all grid sizes. The complex topology and geometry of the panda pose significant challenges to a numerical solver.
@@
The cost of generating cut cells is very much dominated by that of determining the linear system (30), which holds even on the finest grid. Consequently, the cost of the entire initial setup displays a roughly linear growth as the grid size $h$ is reduced. Being the most expensive component of V-cycles and FMG cycles, block smoothing consumes more CPU time than the initial setup on the finest grid, since the $O(h^{-2})$ growth of its cost is higher than the linear growth.
@@ 6.5 FMG efficiency
The panda is a good representative of complex domains while the rotated square in Figure 6(b) that of the other extreme of irregular but simple domains. For the rotated square in Figure 6(b) with $h=\frac{1}{256},\frac{1}{512},\frac{1}{1024}$, we record computational costs (not shown) of the main components. The cost of generating cut cells grows quadratically with respect to the reduction of $h$ while those of all other steps in the initial setup grows linearly; the $O(h^{-2})$ complexity of FMG cycles are also confirmed. In addition, the consumed CPU time in seconds for $h=\frac{1}{1024}$ is 1.89, 2.62, 3.57, and 5.31 for the determination of (30), the entire initial setup, the block smoothing, and all FMG cycles, respectively. The cost ratio of block smoothing over all FMG cycles is $\frac{3.57}{5.31}\approx 0.67$ for the rotated square, which is very close to that ($\frac{2.38}{3.63}\approx 0.66$) for the panda, cf. Table 6. Due to the simple geometry of the rotated square, the cost ratio of the entire setup over all FMG cycles, $\frac{2.62}{5.31}\approx 0.49$, is much smaller than that ($\frac{4.55}{3.63}\approx 1.25$) for the panda test. We sum up main conclusions of the above discussions as follows.
@@
  • Block smoothing consumes $\frac{2}{3}$ of the entire CPU time of FMG cycles and is asymptotically the most expensive component of the proposed method.
(a) Section 6.1 with $h=\frac{1}{512}$
(b) Section 6.4 with $h=\frac{1}{2048}$
Figure 10: Performance of FMG cycles of our method with $\nu_{1}=\nu_{2}=3$ for tests on the unit square $(0,1)^{2}$, the rotated square $\Omega_{r}$, and the panda in Sections 6.1 and 6.4. The ordinates are the relative residuals/errors in $L^{\infty}$ norm, the abscissa is the iteration number of FMG cycles. The condition number of an FMG cycle is indicated by the smallest solution error that remains constant under more multigrid iterations.
@@ 7 Conclusions
Acknowledgments. We acknowledge helpful comments from Shaozhen Cao, Lei Pang, and Chenhao Ye, graduate students at the school of mathematical sciences in Zhejiang University.